OpenStudy (anonymous):

find a polynomial function with integral coefficient that has the given numbers as roots: 5, 1, 4, -1

OpenStudy (anonymous):

(x-5)(x-1)(x-4)(x+1)=0

OpenStudy (amistre64):

i tend to find it simpler to just subtract x from the given roots, not that it makes a difference in the end, but to me its easier to form up than swapping signs ;)

OpenStudy (amistre64):

5, 1, 4, -1 5-x, 1-x, 4-x, -1-x y = (5-x)(1-x)(4-x)(-1-x)